OpenStudy (anonymous):

Given b = - 4 and h = -2, what is the equation of the graph if the parent function is y= \[\sqrt{x}\]

OpenStudy (anonymous):

@mathstudent55 heyy

OpenStudy (anonymous):

@mathstudent55 this is one of the more harder ones that really stump me :(

OpenStudy (anonymous):

Given a = -1, h = 4, and k = -2, what is the equation of the graph if the parent function is y= \[\sqrt{x}\] this is my second question

OpenStudy (mathstudent55):

Where do b and h go in this type pf [arent function?

OpenStudy (anonymous):

i dont know anything thats just what the question says

OpenStudy (mathstudent55):

Is it \(y = \sqrt{b(x - h)} +k \) ?

OpenStudy (mathstudent55):

You must have some studying material that explains the form of the parent function.

OpenStudy (anonymous):

im not sure, just try your best and solve it and ill see if its one of the multiple choice answers when you get your answer

OpenStudy (anonymous):

Use the equation \[y = a \sqrt{x-h}+k\] sometimes the equation can also be expressed as \[y=a \sqrt{x-b}+c\] In addition, there's also the equation\[y=a \sqrt{mx-h}+k\] I would probably use the last one where your b will replace m.

OpenStudy (anonymous):

so what is the answer to the first question?

OpenStudy (mathstudent55):

\(y = -4\sqrt{x + 2}\)

OpenStudy (anonymous):

none of those are the multiple choice they all have -1 in front of it accept for 1 answer

OpenStudy (anonymous):

@mathstudent55 Are you sure? I thought that's where a is, not b.

OpenStudy (mathstudent55):

\( y = \sqrt{-4x + 2}\)

OpenStudy (anonymous):

thats not it either i dont think cuz thats not one of the multiple choice answers

OpenStudy (anonymous):

@mathstudent55 Okay that's better XD

OpenStudy (mathstudent55):

No, I am not sure. I mentioned above that I have not seen the discussion where we learn where b and h are. I am just guessing.

OpenStudy (anonymous):

@MeganxCherry What are your choices?

OpenStudy (anonymous):

this is the multiple choice A\[-1\sqrt{x+4}-2\]

OpenStudy (anonymous):

choice B \[-1\sqrt{x-4}-2\]

OpenStudy (anonymous):

choice C y= \[\sqrt{x-4}-2\]

OpenStudy (mathstudent55):

\(y = \sqrt{-4(x + 2)} \) ?

OpenStudy (anonymous):

Choice D \[-1\sqrt{x-4}+2\]

OpenStudy (mathstudent55):

I don't know. Look in your lesson and see where the b and h belong.

OpenStudy (anonymous):

which one is it

OpenStudy (anonymous):

RPguy do you know? :(

OpenStudy (anonymous):

The thing is that it's really weird to have b and h in the same equation... It's not expressed in a way that I'm familiar of. The only expressions are the ones I've posted above.

OpenStudy (anonymous):

damn so you dont know the answer?

OpenStudy (anonymous):

In this scenario it seems that b = h and h = k, which is just really messed up.

OpenStudy (anonymous):

Give me a couple minutes

OpenStudy (anonymous):

Okay<3

OpenStudy (anonymous):

I think you gave me choices for the second question instead of the first. The second question follows the expression: \[y=a \sqrt{x-h}+k\] and literally just plug in the information you have \[y=-1\sqrt{x-4}-2\] Idk about the first question though

OpenStudy (anonymous):

yeah your right i gave you the multiple from the second question by accident. so what would the answer be what letter choice? your such a great help!

OpenStudy (anonymous):

i got the second question down so im gonna give you the multiple choice for the first one now

OpenStudy (anonymous):

thankyou hold on <3

OpenStudy (anonymous):

A \[\sqrt{4x+8}\]

OpenStudy (anonymous):

B \[\sqrt{4x-8}\]

OpenStudy (anonymous):

C \[\sqrt{-4x-8}\]

OpenStudy (anonymous):

D \[\sqrt{-4x+8}\]

OpenStudy (anonymous):

Take your time, thankyou so much!!

OpenStudy (anonymous):

I think it's D. Using @mathstudent55 's equation \[y=a \sqrt{b(x-h)}+k\] we can plug in our given information \[y=\sqrt{-4(x-2)}\] Which further gives us \[y=\sqrt{-4x+8}\]